Use the following data to answer the following questions

Year Real GDP per Capita (2012 Prices)

2013 $ 47719

2014 $48,116

2015 $48,822

2016 $49,184

2017 $50,010


a. What is the percentage change 2013 to 2017?

la

What is the average annual growth 2013 to 2017?


1
Expert's answer
2022-03-03T12:37:00-0500

a) Percentage change = "\frac{(new - original)}{original
}"

= "\frac{(50,010-47719)}{47719}"

= "\frac{2291}{47719}"

=0.0480


b) Average annual growth = "\frac{total GDP}{Number of years}"

= "\frac{(47719+48,116+48,822+49,184+50,010)}{5}"

= "\frac{243,851}{5}"

= $48,770.20